OpenStudy (anonymous):

can u help me 3.2 is 50%of what?

OpenStudy (anonymous):

isnt it 6.4?

OpenStudy (anonymous):

how do u come up with that i need to know how to figure it up

OpenStudy (anonymous):

because 50% is half and 3.2 is half of 6.4

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[\large \frac{50}{100} \times x = 3.2 \implies \frac{1}{2} \times x = 3.2\] Multiply 2 both the sides: \[\large x = 3.2 \times 2 \implies \color{green}{x = 6.4}\]

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

Welcome dear..